# Epimorphisms between finitely generated algebras

### v2

@@ 4. Quasivarieties with a near unanimity term
Now, as $\langle a,1\rangle\notin A$ and $\bm{B}$ is generated by $A\cup\{\langle a,1\rangle\}$, in order to prove that $\bm{A}\leqslant\bm{B}$ is full in $\mathbb{V}(\mathsf{K})$, it suffices to show that every nonidentity congruence of $\bm{B}$ glues $\langle a,1\rangle$ with some element of $\bm{A}$. But this holds by the above display. As $\bm{A}\leqslant\bm{B}$ is full in $\mathbb{V}(\mathsf{K})$, we can apply Proposition 3.10 obtaining that $\bm{A}\leqslant\bm{B}$ is epic in $\mathbb{V}(\mathsf{K})$ provided that the two conditions in Theorem 3.5 fail. If condition (i) holds, the last part of Proposition 3.10 implies that there exists a nonidentity congruence $\theta$ of $\bm{B}$ with $\theta{\upharpoonright}_{A}=\textup{id}_{A}$, a contradiction with the above display. On the other hand, if condition (ii) holds, $\bm{B}$ embeds into an SI member of $\mathbb{V}(\mathsf{K})$. But this is impossible because the length of $\bm{B}$ is strictly larger than the length of $\bm{A}$, and the latter is an upper bound for the length of the SI members of $\mathbb{V}(\mathsf{K})$. ∎
@@
The implicative reducts of Boolean algebras are called implication algebras [1]. These form a variety $\mathsf{IA}$ that is congruence distributive, but lacks a near unanimity term [27, Lem. 3]. We will prove that Proposition 4.6 fails for $\mathsf{IA}$. Let $\bm{B}$ be the implicative reduct of the powerset Boolean algebra $\mathcal{P}(\mathbb{N})$ and $\bm{A}$ the subalgebra of $\bm{B}$ with universe $\mathcal{P}(\mathbb{N})-\{\emptyset\}$. We will show that $\bm{A}\leqslant\bm{B}$ is full in $\mathsf{IA}$, but that $\bm{B}$ cannot be obtained as a subdirect product $\bm{B}\leqslant\bm{A}_{1}\times\dots\times\bm{A}_{n}$ in which each $\bm{A}_{i}$ is FSI. The latter holds because the only FSI implication algebra is the implicative reduct of the two-element Boolean algebra, while $\bm{B}$ is infinite. Then we turn to prove that $\bm{A}\leqslant\bm{B}$ is full in $\mathsf{IA}$. Clearly, $\bm{A}\leqslant\bm{B}$ is proper and almost total. Moreover, as the congruences of $\bm{B}$ coincide with those of the powerset Boolean algebra $\mathcal{P}(\mathbb{N})$, every nonidentity congruence of $\bm{B}$ glues $\emptyset$ with some element of $\bm{A}$ [^6] Although we do not need it here, we remark that $\bm{A}\leqslant\bm{B}$ is indeed fully epic in $\mathsf{IA}$ (cf. [5, Prop. 4.5]).
@@ 5. Congruence permutable varieties
A variety $\mathsf{K}$ is congruence permutable iff $\theta_{1}+^{\bm{A}}\theta_{2}=\theta_{1}\circ\theta_{2}$ for every $\bm{A}\in\mathsf{K}$ and $\theta_{1},\theta_{2}\in\mathsf{Con}(\mathsf{K})$.
Varieties that are both congruence distributive and congruence permutable are called arithmetical. Moreover, a class of algebras is said to be universal when it is closed under $\mathbb{I},\mathbb{S}$, and $\mathbb{P}_{\!\textsc{{u}}}$ or, equivalently, when it can be axiomatized by universal sentences (see, e.g., [7, Thm. V.2.20]). The task of determining whether an arithmetical variety, whose class of FSI members is universal, has the ES property can be simplified as follows [8, Thm. 22]:
@@
The implication from left to right follows from Proposition 3.2. To prove the other implication, we reason by contraposition. Suppose that $\mathsf{K}$ lacks the weak ES property. By Corollary 3.8 there exist a finitely generated $\bm{B}\in\mathsf{K}$ and $\bm{A}\leqslant\bm{B}$ fully epic in $\mathsf{K}$. By the next proposition (which will be established in the rest of the section) we conclude that $\bm{B}\in\mathsf{K}_{\textsc{fsi}}$. ∎
@@
where the first equality follows from the fact that $\textup{Cg}^{\bm{B}}(\theta_{1}{\upharpoonright}_{A}+^{\bm{A}}\theta_{2}{\upharpoonright}_{A})$ and $\textup{Cg}^{\bm{B}}(\theta_{1}{\upharpoonright}_{A})+^{\bm{B}}\textup{Cg}^{\bm{B}}(\theta_{2}{\upharpoonright}_{A})$ coincide as they are both the smallest congruence of $\bm{B}$ containing $\theta_{1}{\upharpoonright}_{A}$ and $\theta_{2}{\upharpoonright}_{A}$, and the second equality is a consequence of Lemma 5.5. ∎
@@ 6. Arithmeticity is often necessary
We recall that quasivarieties contain free algebras. More precisely, for every quasivariety $\mathsf{K}$ and set $X$ there exists an algebra $\bm{T}_{\mathsf{K}}(X)$ that is free in $\mathsf{K}$ over $X$ (see, e.g., [15, Prop. 2.1.10]). When $X=\{x,y\}$, we will write $\bm{T}_{\mathsf{K}}(x,y)$ instead of $\bm{T}_{\mathsf{K}}(X)$.
@@
Recall that $\bm{A}/\theta\leqslant\bm{B}/\theta$ is full in $\mathsf{K}$. On the other hand, as $\mathsf{K}$ has the weak ES property by assumption, $\bm{A}/\theta\leqslant\bm{B}/\theta$ is not epic in $\mathsf{K}$. Therefore, we can apply Corollary 3.11 obtaining that $\bm{A}/\theta\in\mathbb{S}(\mathsf{K}_{\textsc{rfsi}})$. Since $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras by assumption, $\bm{A}/\theta$ is trivial or RFSI. Thus, $\theta{\upharpoonright}_{A}=A\times A$ or $\theta{\upharpoonright}_{A}\in\mathsf{Irr}_{\mathsf{K}}(\bm{A})$, and hence the above display implies that
@@
  2. (ii)
    $\mathsf{DL}$ is congruence distributive and $\mathsf{DL}_{\textsc{fsi}}$ is closed under nontrivial subalgebras;
@@
    $\mathsf{SL}$ has the ES property and $\mathsf{SL}_{\textsc{fsi}}$ is closed under nontrivial subalgebras. ∎
@@
A variety $\mathsf{K}$ is said to be filtral [26] when for every subdirect product $\bm{A}\leqslant\prod_{i\in I}\bm{A}_{i}$ where $\{\bm{A}_{i}:i\in I\}$ is a family of SI members of $\mathsf{K}$ and every $\theta\in\mathsf{Con}(\bm{A})$ there exists a filter $F$ on $I$ such that
@@
where $\llbracket a=b\rrbracket\coloneqq\{i\in I:a(i)=b(i)\}$. Moreover, a variety $\mathsf{K}$ is discriminator [32] when $\mathsf{K}=\mathbb{V}(\mathsf{M})$ for some class of algebras $\mathsf{M}$ for which there exists a term $\varphi(x,y,z)$ such that for every $\bm{A}\in\mathsf{M}$ and $a,b,c\in A$,
@@ Acknowledgements.
The second author was supported by the ayuda PREP2022-000927 financiada por MICIU/AEI/$10$.13039130391303913039/$501100011033$ y por FSE+ as part of the proyecto PID2022202220222022-$141529$NB-C21212121 de investigación financiado por MICIU/AEI/$10$.13039130391303913039/$501100011033$ y por FEDER, UE.

### v3

@@ 4. Quasivarieties with a near unanimity term
Now, as $\langle a,1\rangle\notin A$ and $\bm{B}$ is generated by $A\cup\{\langle a,1\rangle\}$, in order to prove that $\bm{A}\leqslant\bm{B}$ is full in $\mathbb{V}(\mathsf{K})$, it suffices to show that every nonidentity congruence of $\bm{B}$ glues $\langle a,1\rangle$ with some element of $\bm{A}$. But this holds by the above display. As $\bm{A}\leqslant\bm{B}$ is full in $\mathbb{V}(\mathsf{K})$, we can apply Proposition 3.10 obtaining that $\bm{A}\leqslant\bm{B}$ is epic in $\mathbb{V}(\mathsf{K})$ provided that the two conditions in Theorem 3.5 fail. If condition (i) holds, the last part of Proposition 3.10 implies that there exists a nonidentity congruence $\theta$ of $\bm{B}$ with $\theta{\upharpoonright}_{A}=\textup{id}_{A}$, a contradiction with the above display. On the other hand, if condition (ii) holds, $\bm{B}$ embeds into an SI member of $\mathbb{V}(\mathsf{K})$. But this is impossible because the length of $\bm{B}$ is strictly larger than the length of $\bm{A}_{1}$, and the latter is an upper bound for the length of the SI members of $\mathbb{V}(\mathsf{K})$. ∎
@@
The implicative subreducts of Boolean algebras are called implication algebras [1]. These form a variety $\mathsf{IA}$ that is congruence distributive, but lacks a near unanimity term [27, Lem. 3]. We will prove that Proposition 4.6 fails for $\mathsf{IA}$. Let $\bm{B}$ be the implicative reduct of the powerset Boolean algebra $\mathcal{P}(\mathbb{N})$ and $\bm{A}$ the subalgebra of $\bm{B}$ with universe $\mathcal{P}(\mathbb{N})-\{\emptyset\}$. We will show that $\bm{A}\leqslant\bm{B}$ is full in $\mathsf{IA}$, but that $\bm{B}$ cannot be obtained as a subdirect product $\bm{B}\leqslant\bm{A}_{1}\times\dots\times\bm{A}_{n}$ in which each $\bm{A}_{i}$ is FSI. The latter holds because the only FSI implication algebra is the implicative reduct of the two-element Boolean algebra, while $\bm{B}$ is infinite. Then we prove that $\bm{A}\leqslant\bm{B}$ is full in $\mathsf{IA}$. Clearly, $\bm{A}\leqslant\bm{B}$ is proper and almost total. Moreover, as the congruences of $\bm{B}$ coincide with those of the powerset Boolean algebra $\mathcal{P}(\mathbb{N})$, every nonidentity congruence of $\bm{B}$ glues $\emptyset$ with some element of $\bm{A}$ [^6] Although we do not need it here, we remark that $\bm{A}\leqslant\bm{B}$ is indeed fully epic in $\mathsf{IA}$ (cf. [5, Prop. 4.5]).
@@ 5. Congruence permutable varieties
A variety $\mathsf{K}$ is congruence permutable iff $\theta_{1}+^{\bm{A}}\theta_{2}=\theta_{1}\circ\theta_{2}$ for every $\bm{A}\in\mathsf{K}$ and $\theta_{1},\theta_{2}\in\mathsf{Con}(\bm{A})$.
Varieties that are both congruence distributive and congruence permutable are called arithmetical. Moreover, a class of algebras is said to be universal when it is closed under $\mathbb{I},\mathbb{S}$, and $\mathbb{P}_{\!\textsc{{u}}}$ or, equivalently, when it can be axiomatized by universal sentences (see, e.g., [7, Thm. V.2.20]). The task of determining whether an arithmetical variety, whose class of FSI members is universal, has the ES property can be simplified as follows [8, Thm. 6.8]:
@@
The implication from left to right follows from Proposition 3.2. To prove the other implication, we reason by contraposition. Suppose that $\mathsf{K}$ lacks the weak ES property. By Corollary 3.8 there exist a finitely generated $\bm{B}\in\mathsf{K}$ and $\bm{A}\leqslant\bm{B}$ fully epic in $\mathsf{K}$. Using the next proposition (which will be established in the rest of the section) we conclude that $\bm{B}\in\mathsf{K}_{\textsc{fsi}}$. ∎
@@
where the first equality follows from the fact that $\textup{Cg}^{\bm{B}}(\theta_{1}{\upharpoonright}_{A}+^{\bm{A}}\theta_{2}{\upharpoonright}_{A})$ and $\textup{Cg}^{\bm{B}}(\theta_{1}{\upharpoonright}_{A})+^{\bm{B}}\textup{Cg}^{\bm{B}}(\theta_{2}{\upharpoonright}_{A})$ are both the smallest congruence of $\bm{B}$ containing $\theta_{1}{\upharpoonright}_{A}$ and $\theta_{2}{\upharpoonright}_{A}$, and the second equality is a consequence of Lemma 5.5. ∎
@@ 6. Arithmeticity is often necessary
We recall that quasivarieties contain free algebras. More precisely, for every quasivariety $\mathsf{K}$ and nonempty set $X$ there exists an algebra $\bm{T}_{\mathsf{K}}(X)$ that is free in $\mathsf{K}$ over $X$ (see, e.g., [15, Prop. 2.1.10]). When $X=\{x,y\}$, we will write $\bm{T}_{\mathsf{K}}(x,y)$ instead of $\bm{T}_{\mathsf{K}}(X)$.
@@
Recall that $\bm{A}/\theta\leqslant\bm{B}/\theta$ is full in $\mathsf{K}$. On the other hand, as $\mathsf{K}$ has the weak ES property by assumption, $\bm{A}/\theta\leqslant\bm{B}/\theta$ is not epic in $\mathsf{K}$. Therefore, we can apply Corollary 3.11 obtaining that $\bm{A}/\theta\in\mathbb{S}(\mathsf{K}_{\textsc{rsi}})\subseteq\mathbb{S}(\mathsf{K}_{\textsc{rfsi}})$. Since $\mathsf{K}_{\textsc{rfsi}}$ is closed under nontrivial subalgebras by assumption, $\bm{A}/\theta$ is trivial or RFSI. Thus, $\theta{\upharpoonright}_{A}=A\times A$ or $\theta{\upharpoonright}_{A}\in\mathsf{Irr}_{\mathsf{K}}(\bm{A})$, and hence the above display implies that
@@
    $\mathsf{SL}$ has the ES property and $\mathsf{SL}_{\textsc{fsi}}$ is closed under nontrivial subalgebras;
  2. (ii)
@@
    $\mathsf{DL}$ is congruence distributive and $\mathsf{DL}_{\textsc{fsi}}$ is closed under nontrivial subalgebras. ∎
@@
A variety $\mathsf{K}$ is said to be filtral [26] when for every subdirect product $\bm{A}\leqslant\prod_{i\in I}\bm{A}_{i}$ where $\{\bm{A}_{i}:i\in I\}$ is a family of SI members of $\mathsf{K}$ and every $\theta\in\mathsf{Con}(\bm{A})$ there exists a filter $F$ over $I$ such that
@@
where $\llbracket a=b\rrbracket\coloneqq\{i\in I:a(i)=b(i)\}$. Moreover, a variety $\mathsf{K}$ is a discriminator variety [31] when $\mathsf{K}=\mathbb{V}(\mathsf{M})$ for some class of algebras $\mathsf{M}$ for which there exists a term $\varphi(x,y,z)$ such that for every $\bm{A}\in\mathsf{M}$ and $a,b,c\in A$,
@@ Acknowledgements.
The second author was supported by the ayuda PREP2022202220222022-$000927$ financiada por MICIU/AEI/$10$.13039130391303913039/$501100011033$ y por FSE+ as part of the proyecto PID2022202220222022-$141529$NB-C21212121 de investigación financiado por MICIU/AEI/$10$.13039130391303913039/$501100011033$ y por FEDER, UE.